Can someone help me solve this equation please?
9x^2=14
Thank you.

Answer by stanbon(75887)

9x^2=14
x^2=14/9
x-sqrt14/9) or x=-sqrt(14/9)
Multiply numerator and denominator by sqrt9 to get:
x=[sqrt126]/9 or x=-[sqrt126]/9
Cheers,
Stan H.